Graph the ellipse, then choose the correct foci.
x^2 + 25y^2 = 100

Answer by fractalier(6550)

First, put it in standard form by dividing by 100...we get
x^2 + 25y^2 = 100
x^2/100 + y^2/4 = 1
Here a^2 = 100 and b^2 = 4
We find the focal distance via
c^2 = a^2 - b^2
c^2 = 100 - 4 = 96 so that
c = sqrt(96) = 4*sqrt(6) and the foci are then
(-4√6, 0) and (4√6, 0)